Resolution for the issue, for the archives.
Power cycling the 3573 library did clear the Drive 4 error.
Volume 001005 was able to be moved/unloaded to a slot.
Media attention light remained on.
Moving 001005 from IO slot to I/O door caused IO Station Bad Tape indicator to appear.
Removed 001005 from tape library, 3573, cleared both Media Attention and Bad Tape light.
Reloaded 001005 back into library, Drive 4, DSPTAP of 001005, *SAVRST, worked fine with no error.
Currently running another DUP, different volume as source in Drive 4, with no issue.
Volume 001005 usable, but still in *ERR status within BRMS.
Beginning with V7R1 9=Remove volume error status
Ran this on 001005.
At this point, we don't have solid confirmation of cause of the failure drive 4 or volume 001005.
It was many moons ago, but I remember a similar issue, neither the drive nor the volume was replaced.
I reran the DUPMEDBRM on the same Drive4 with same input volume, 001005, everything ran fine, no errors or failures.
This confirmed that neither Drive4 nor volume 001005 was the cause of the error.
After sending library service logs and drive logs, and requesting a PE review the logs, here is IBM's initial response.
There is new LTO5 HH FC drive firmware, (not yet released) that should resolve the issue.
Current LTO5 HH FC firmware is D8D5, D8D9 available only when recommended by PE.
Summarizing, this is the 4th incident in the last 2 1/2 years, while on Power7 and/or LTO5, where NEWER firmware was recommended by IBM for the resolution.
In all cases, Power7 (VPD issues) or LTO5, critical errors could have been avoided if newer firmware would have been installed.
Any feedback from the group on firmware related errors, has anyone experienced any of these.
Note: If there was a scheduled restore process from LPAR A to LPAR B that night, which needed volume 001005, this process would have failed with no immediate resolution.
From: MIDRANGE-L [mailto:midrange-l-bounces@xxxxxxxxxxxx] On Behalf Of Steinmetz, Paul
Sent: Monday, May 12, 2014 10:14 AM
To: 'Midrange Systems Technical Discussion'
Subject: BRMS DUP failure CPF5349 , CPP630C - 3573 LTO5 HH FC Drvie failure
Following a successful BRMS full save on LPAR A, BRMS auto DUP job experienced a CPF5349, CPP630C on the source volume 001005.
Vol 001005 was used for successful restore on LPAR B, 2 hour prior to the failure on LPAR A.
BRMS marked 001005 in an *ERR status, normal.
LTO 5 HH FC Drive 4 is in error status.
Library warning trace logs show a Drive Warn or Crit Tape Alert flag.
001005 is currently stuck in Drive 4 of the 3573, cannot be unloaded either through BRMS or 3573 GUI move media commands.
Not sure of the cause, bad volume 001005 or bad drive 4.
3573 operator guide suggests to power cycle the library.
Our previous 3582 library had a button on the drive itself that you could reach within the library and force remove a stuck volume.
I'm not finding this same functionality with the 3573.
Any suggestions on proper next procedures?
IBM i Systems Administrator
Pencor Services, Inc.
462 Delaware Ave
Palmerton Pa 18071